OpenStudy (anonymous):

find the value(s) of x for which f(x)=g(x): f(x)=x^2+2x+1and g(x)=3x+3

OpenStudy (anonymous):

set the two equations equal to each other and solve for x so; x^2+2x+1 = 3x+3

OpenStudy (anonymous):

I know but I don't understand this is just an example so I can complete the real question

OpenStudy (loser66):

so, post the real question to get help!

OpenStudy (anonymous):

its close to that if you solve that then i got my answer

OpenStudy (jazzyfa30):

\[f(x)+x^2+2x+1=3x+3\]

OpenStudy (loser66):

f(x)+...?

OpenStudy (anonymous):

find the value(s) of x for which f(x)=g(x): f(x)=x^2+2x+1and g(x)=3x+3

OpenStudy (loser66):

from robz8 you have x^2 +2x +1 = 3x+3 x^2+2x+1-3x-3=0 (x-2)(x+1)=0 --> x=2 and x= -1